OpenStudy (jjuden):

select the correct numbers to create an equation that is true for all values of x 2(4x+3)(3x+5)=Ax^2+58x+B a A=6, B=8 b A=12, B=15 c A=24, B=30 d A=29, B=58

OpenStudy (astrophysics):

First distribute (multiply), (4x+3) over (3x+5)

OpenStudy (astrophysics):

Once we distribute it all we can equate the two sides

OpenStudy (astrophysics):

Good, now distribute the 2 over it

OpenStudy (jjuden):

24x^2+58x+30

OpenStudy (astrophysics):

Perfect, so we have \[24x^2+58x+30=Ax^2+58x+B\]

OpenStudy (astrophysics):

Now we can equate the two sides, notice \[24x^2=Ax^2\]\[58x=58x\]\[30=B\]

OpenStudy (astrophysics):

Haha, well I gave away B, but note the coefficient of A, which would be?

OpenStudy (jjuden):

so a=24?

OpenStudy (astrophysics):

Yes

OpenStudy (jjuden):

thank you